OpenStudy (anonymous):

The derivative of f(x) is defined by f'(x) =lim(h>o) (f(x+h) -f(x)/ h). Use this definition to find f'(x) for f(x)=x^3 +2x

OpenStudy (paxpolaris):

\[\lim_{h \to 0} {\left[ \left( x+h \right)^3+2\left( x+h \right) \right]-\left[ x^3+2x \right]\over h}\]

OpenStudy (paxpolaris):

go for it. simplify